OpenStudy (anonymous):

I don't understand why 8 is the answer for this problem? Using this-(x - 2y)(x + 2y) = 4 A. x^2-4y^2 B. 8 (which is greater)

OpenStudy (anonymous):

Foil it out.

OpenStudy (bahrom7893):

(x - 2y)(x + 2y) = x^2-4y^2 = 4

OpenStudy (bahrom7893):

4<8 therefore 8 is larger!

OpenStudy (anonymous):

ok you get x^2+4y=4

OpenStudy (bahrom7893):

not +, -

OpenStudy (anonymous):

Is the - after the word 'Using' part of the equation?

OpenStudy (bahrom7893):

and that - in front is not supposed to be there polpak..

OpenStudy (anonymous):

is that a minus sign in the front? not that it makes any difference

OpenStudy (anonymous):

people need to learn to properly use colons.

OpenStudy (anonymous):

i meant x^2-4y^2

OpenStudy (bahrom7893):

lol yea it doesn

OpenStudy (bahrom7893):

-4 is still smaller than 8

OpenStudy (anonymous):

no it does not have a minus infront